75.4 g of Na2S, sodium sulphide is mixed with water to form 4.0 L of solution. What is the concentration of the sodium ions, Na+1 in the solution? Hint: write out dissociation reaction


1
Expert's answer
2021-03-25T04:09:26-0400

Molar concentration = (mass/volume) ÷ molar mass

= (75.4/4) ÷ 78

= 0.242M


Na2S = 2Na + S

Concentration of Na = 2 × 0.242M = 0.484M